# SparseLIF: High-Performance Sparse LiDAR-Camera Fusion for 3D Object Detection

### v1

@@ Abstract
Sparse 3D detectors have received significant attention since the query-based paradigm embraces low latency without explicit dense BEV feature construction. However, these detectors achieve worse performance than their dense counterparts. In this paper, we find the key to bridging the performance gap is to enhance the awareness of rich representations in two modalities. Here, we present a high-performance fully sparse detector for end-to-end multi-modality 3D object detection. The detector, termed SparseLIF, contains three key designs, which are (1) Perspective-Aware Query Generation (PAQG) to generate high-quality 3D queries with perspective priors, (2) RoI-Aware Sampling (RIAS) to further refine prior queries by sampling RoI features from each modality, (3) Uncertainty-Aware Fusion (UAF) to precisely quantify the uncertainty of each sensor modality and adaptively conduct final multi-modality fusion, thus achieving great robustness against sensor noises. By the time of submission (2024/03/08), SparseLIF achieves state-of-the-art performance on the nuScenes dataset, ranking 1st on both validation set and test benchmark, outperforming all state-of-the-art 3D object detectors by a notable margin. The source code will be released upon acceptance.
@@ 1 Introduction
LiDAR-camera-based 3D detection is essential for accurate and robust autonomous driving systems. The two modalities naturally provide complementary information, i.e., the camera offers high-resolution semantic information while LiDAR provides accurate geometric information. Therefore, camera and LiDAR sensors have been simultaneously deployed for reliable recognition and localization of objects in 3D space.
@@
This paper presents SparseLIF, a high-performance fully sparse multi-modality 3D object detector that outperforms all other dense counterparts and sparse detectors. SparseLIF bridges the performance gap by enhancing the awareness of rich LiDAR and camera representations in three aspects, i.e., query generation, feature sampling and multi-modality fusion. First, we argue that the convention [68], which randomly generates queries, will suffer from extra efforts in learning to move the query proposals towards ground-truth targets. Here, we propose the Perspective-Aware Query Generation (PAQG) module to ease learning. In particular, PAQG injects a lightweight perspective detector composed of the coupled 2D and monocular-3D sub-networks on image features to predict and transform top-scored 3D proposals into query proposals. These input-dependent proposals will narrow the learning path toward ground-truth targets, thus enhancing the awareness of rich contexts in high-resolution images. Second, these queries with perspective priors will interact with features from two modalities via the RoI-Aware Sampling (RIAS) module. Instead of resorting to cumbersome global attention, the module locates the region of interest and then sample complementary features at merely several reference points under the guidance of prior queries, thus conforming to the fully sparse paradigm and enjoying low latency. Third, we observe that in realistic scenarios, LiDAR and camera usually suffer from various sensor problems as shown in Fig. 3, which will make sensor inputs unreliable and uncertain, thus degrading the performance of multi-modality detectors. Hence, we propose the Uncertainty-Aware Fusion (UAF) module to precisely quantify the uncertainty of each modality and guide our model to focus on the trustworthy modality in multi-modality fusion, thus achieving great robustness against sensor noises. Finally, our contributions and achievements are summarized as follows.
@@ 2.3 LiDAR-Camera-based 3D Object Detection
Recent works utilize the sparse query-based paradigm without explicit view transformation. Transfusion obtains object queries from LiDAR points and then fuses queries with rich image features using a transformer block. CMT [71] further develops an end-to-end feature interaction framework for multi-modality fusion. UniTR [64] introduces a modality-agnostic transformer encoder to proceed with unified modeling and shared parameters. Despite great success, the expensive global attention buries the advantages of the sparse paradigm and makes it difficult to benefit from long-term temporal information.
Another stream of works explores the fully sparse paradigm. SparseFusion [83] poses detectors on each modality and fuses features of detected instances. However, the two-stage paradigm suffers from the limited performance of modality-specific detectors. FUTR3D [7] generalizes the fully sparse paradigm by initializing 3D reference points and projecting them into all available modalities to sample features. Although methods have recently achieved good performance, there remains a notable performance gap compared to dense counterparts.
@@ 3.1 Perspective-Aware Query Generation
Figure 2: Motivations and details of our proposed PAQG module. (a) 3D detectors struggle with low sensitivity when detecting distant and small objects. (b) 2D detectors demonstrate excellent pixel-wise perception capabilities on such objects. (c) the PAQG module adopts the coupled 2D and monocular-3D sub-networks to predict dense boxes under the supervision of a perspective loss. Then, we pick top-ranked boxes to propose high-quality queries, and then interact with camera features via a cross-attention module.
The adopted perspective detector in the PAQG module consists of the coupled 2D (e.g. FCOS [61]) and monocular-3D (e.g. FCOS3D [66]) sub-networks. Taken the multi-view/scale image features $X_{\text{cam}}$ as input, the monocular-3D sub-network predicts raw 3D attributes, i.e., depths $\mathbf{d}$, rotation angles, sizes, and velocities throughout different views. Simultaneously, the 2D sub-network predicts corresponding 2D attributes, i.e., center coordinates $[\mathbf{c_{x}},\mathbf{c_{y}}]$, confidence scores, and category labels. For each view $v$, we project the box centers into 3D space based on corresponding camera extrinsic ${E_{v}}$ and intrinsic ${I_{v}}$, i.e.,
@@
The 3D center $\mathbf{c}^{3D}$ will combine with the predicted size, rotation angle, and velocity to form 3D boxes. Then, we perform non-maximum suppression in 3D space to filter intersecting boxes and pick the top $N_{k}$ boxes ranked by confidence scores to initialize queries via a cross-attention module. Formally,
@@
where $\mathcal{P}_{\text{cam}}^{v}(c^{3D}_{i})$ projects the 3D center $c^{3D}_{i}$ to $v$-th image using corresponding camera parameters. Besides, $\mathcal{V}$ denotes the set of hit views. $\mathcal{BS(\cdot)}$ denotes the bilinear sampling function. Since some objects may be overlooked, we preserve $N_{r}$ randomly initialized query boxes. Finally, our PAQG module generates total $N_{q}=N_{k}+N_{r}$ query proposals. This way, our PAQG module provides input-dependent query proposals to elevate the understanding of perspective priors for 3D detectors, thereby aiding in detecting distant and small objects.
@@ 3.2.3 Channel-Spatial Correlation Aware Mixing
where $W_{s}$ is the spatial correlation shared across different channels.
After channel-spatial correlation aware mixing, the features are flattened and aggregated by a linear layer. The final regression and classification predictions are computed by two MLPs, respectively.
@@ 3.3 Uncertainty-Aware Fusion
$\bar{q}_{i}=f_{UA}(F_{\text{cam}},U_{\text{cam}},F_{\text{lid}},U_{\text{lid}}),$ (9)
where $\bar{q}_{i}$ and $f_{UA}$ are the refined query feature and uncertainty-aware fusion function, respectively. Besides, $U_{\text{cam}}$ and $U_{\text{lid}}$ are the uncertainty of two modalities.
@@
where $f_{\text{dist}}$ is the distance predictor consisted with MLPs.
@@
[…]
@@
${}^{\ddagger}$ [41] LC 24.5 23.3 30.8 23.3 13.1 71.4 74.2
${}^{\ddagger}$ [16] LC 24.3 23.8 34.5 32.8 13.3 75.5 74.9
${}^{\ddagger}$ [44] LC 24.2 22.7 32.0 22.2 13.0 75.0 76.1
${}^{\ddagger}$ [75] LC 23.5 23.3 32.8 22.6 13.0 75.6 76.3
${}^{\ddagger}$ [71] LC 23.3 22.0 27.1 21.2 12.7 75.3 77.0
${}^{\ddagger}$ [4] LCT 22.9 22.9 30.2 22.5 13.5 76.8 77.2
${}^{\ddagger}$ [20] LC 23.4 22.8 27.8 20.4 12.4 76.6 77.6
[…]
@@
#1 LC V2-99 VoxelNet 69.8 73.3
#2 LC V2-99 VoxelNet 71.0 74.3
#3 LC V2-99 VoxelNet 70.5 74.1
#4 LC V2-99 VoxelNet 71.2 74.6
Table 4: Performance analysis of our PAQG module on detection distances and small object classes, based on SparseLIF-S. The $AP$ scores of traffic cone (T.C.) and barrier at $30m$- are missing since corresponding annotations are unavailable.
PAQG
00-$10m$
$10$-$20m$
$20$-$30m$
$30m$-
mAP
75.1
74.3
65.5
56.9
mAP
76.2 74.8 66.8 58.5
T.C.
82.5
80.6
69.1
T.C.
83.8 82.7 70.4
Barrier
72.1
77.5
54.5
Barrier
76.4 80.0 63.1
@@ 4.1.1 Implementation Details
Each model is trained end-to-end using the AdamW optimizer on eight NVIDIA Tesla-A100 GPUs with a total batch size of $8$. For fair comparisons, we apply the query-denoising strategy [30], commonly used in sparse detection heads, to address the unstable matching problem. Each model is trained for $24$ epochs with a learning rate of $42e-$ 4.
@@ 4.2 Comparisons with State-of-the-Art 3D Object Detectors
As shown in the top part of Tab. 1, without using any test-time augmentation (TTA) or model ensemble, our SparseLIF-T achieves state-of-the-art single-model performance, reaching $77.0\%$ $NDS$ on the nuScenes test benchmark, significantly outperforming all other 3D detectors by a notable margin. In particular, we outperform the most competitive method FusionFormer [19] by $1.9\%$ $NDS$ without using any external training data. Regarding the test benchmark leaderboard in the bottom part of Tab. 1, many competitive methods [41, 16, 44, 75, 71, 4, 20] adopt very complex model ensemble (e.g. assembling models with different voxel sizes, BEV sizes, backbones/FPNs/heads) and TTA, to strive for top ranking on the test leaderboard, e.g. the previous No.1 method EA-LSS [20] improves by $3.2\%$ $NDS$ than its single model in this way. Contrarily, we only use very simple self-model ensemble without TTA for SparseLIF-T (i.e., $0.7\%$ $NDS$ improvement), achieving the best performance of $77.7\%$ $NDS$ and ranking 1st on the test leaderboard by the time of submission.
@@
We also conduct latency analysis on the nuScenes dataset. We implement SparseLIF using Pytorch without any acceleration operations. The overall latency of SparseLIF-S is $340ms$, while the overall latency of BEVFusion [34] is $1610ms$, $4.7$ times slower than ours on the same device. In detail, the detection head (including the PAQG module, the RIAS module, and the UAF module, etc.) only takes about $40ms$ ($\approx 11.8\%$), while the camera and LiDAR backbones take the rest of the time, which demonstrates the efficiency of our detector. We can further speed up our detector by configuring the backbones.
Table 5: Robustness studies of our multi-modality detector under challenging scenarios: LiDAR malfunction, camera malfunction and unsynchronization. Abbreviations: construction vehicle (C.V.), pedestrian (Ped.), and traffic cone (T.C.).
@@ 4.3 Ablation Studies
In Tab. 3, we conduct ablation studies on the nuScenes validation set to evaluate the key components in our multi-modality framework, based on the state-of-the-art single-frame detector SparseLIF-S, which yields highly convincing proofs. Adopting the PAQG module to generate high-quality query proposals, the $mAP$ and $NDS$ are improved by $1.2\%$ and $1.0\%$ respectively. Adopting the UAF module to conduct multi-modality fusion, the $mAP$ and $NDS$ are improved by $0.7\%$ and $0.8\%$ respectively. When SparseLIF-S assembles both modules, the best performance is reached: $71.2\%$ $mAP$ and $74.6\%$ $NDS$.
@@ 4.4 Robustness Studies
We further visualize the predictions of our SparseLIF under the most challenging scenario of limited LiDAR FoV angle of $120^{\circ}$. As presented in Fig. 4, our SparseLIF precisely detects objects in golden circles with malfunctioning LiDAR input, showing the remarkable robustness of our multi-modality detector attributed to the proposed UAF module.
@@ 5 Conclusion
This paper proposes a high-performance fully sparse detector termed SparseLIF for LiDAR-camera-based 3D object detection. The proposed detector achieves state-of-the-art by enhancing the awareness of rich representations in two modalities. In particular, SparseLIF consists of (1) the PAQG module, which generates high-quality 3D queries with perspective priors to facilitate the perception of small and distant objects; (2) the RIAS module, which further refines prior queries by RoI feature sampling to embrace the fully sparse paradigm with the capability of low latency and integration of more temporal frames; (3) the UAF module, which quantifies the uncertainty of each modality for multi-modality fusion to enhance perception robustness against sensor noises. The experimental results demonstrate the superiority of our proposed method over all state-of-the-art 3D object detectors on the nuScenes benchmark. In the future, we will explore applications of SparseLIF on other tasks, such as occupancy prediction.

### v2

@@ Abstract
Sparse 3D detectors have received significant attention since the query-based paradigm embraces low latency without explicit dense BEV feature construction. However, these detectors achieve worse performance than their dense counterparts. In this paper, we find the key to bridging the performance gap is to enhance the awareness of rich representations in two modalities. Here, we present a high-performance fully sparse detector for end-to-end multi-modality 3D object detection. The detector, termed SparseLIF, contains three key designs, which are (1) Perspective-Aware Query Generation (PAQG) to generate high-quality 3D queries with perspective priors, (2) RoI-Aware Sampling (RIAS) to further refine prior queries by sampling RoI features from each modality, (3) Uncertainty-Aware Fusion (UAF) to precisely quantify the uncertainty of each sensor modality and adaptively conduct final multi-modality fusion, thus achieving great robustness against sensor noises. By the time of paper submission, SparseLIF achieves state-of-the-art performance on the nuScenes dataset, ranking 1st on both validation set and test benchmark, outperforming all state-of-the-art 3D object detectors by a notable margin.
@@ 1 Introduction
LiDAR-camera-based 3D detection is essential for accurate and robust autonomous driving systems. The two modalities naturally provide complementary information, i.e., the camera offers high-resolution semantic information while LiDAR provides accurate geometric information. Therefore, camera and LiDAR sensors have been simultaneously deployed for reliable 3D object detection.
@@
This paper presents SparseLIF, a high-performance fully sparse multi-modality 3D object detector that outperforms all other dense counterparts and sparse detectors. SparseLIF bridges the performance gap by enhancing the awareness of rich LiDAR and camera representations in three aspects, i.e., query generation, feature sampling and multi-modality fusion. First, we argue that the convention [68], which randomly generates queries, will suffer from extra efforts in learning to move the query proposals towards ground-truth targets. Here, we propose the Perspective-Aware Query Generation (PAQG) module to ease learning. In particular, PAQG injects a lightweight perspective detector composed of the coupled 2D and monocular-3D sub-networks on image features to predict and transform top-scored 3D proposals into query proposals. These input-dependent proposals will narrow the learning path toward ground-truth targets, thus enhancing the awareness of rich contexts in high-resolution images. Second, these queries with perspective priors will interact with features from two modalities via the RoI-Aware Sampling (RIAS) module. Instead of resorting to cumbersome global attention, the module locates the region of interest and then sample complementary features at merely several reference points under the guidance of prior queries, thus conforming to the fully sparse paradigm and enjoying low latency. Third, we observe that in realistic scenarios, LiDAR and camera usually suffer from various sensor problems as shown in Fig. 3, which will make sensor inputs unreliable and uncertain, thus degrading the performance of multi-modality detectors. Hence, we propose the Uncertainty-Aware Fusion (UAF) module to precisely quantify the uncertainty of each modality and guide our model to focus on the trustworthy modality in multi-modality fusion, thus achieving great robustness against sensor noises. Our contributions are summarized as follows.
@@ 2.3 LiDAR-Camera-based 3D Object Detection
Recent works utilize the sparse query-based paradigm without explicit view transformation. Transfusion [61] obtains object queries from LiDAR points and then fuses queries with rich image features using a transformer block. CMT [71] further develops an end-to-end feature interaction framework for multi-modality fusion. UniTR [64] introduces a modality-agnostic transformer encoder to proceed with unified modeling and shared parameters. Despite great success, the expensive global attention buries the advantages of the sparse paradigm and makes it difficult to benefit from long-term temporal information.
Another stream of works explores the fully sparse paradigm. SparseFusion [83] poses detectors on each modality and fuses features of detected instances. However, the two-stage paradigm suffers from the limited performance of modality-specific detectors. FUTR3D [7] generalizes the fully sparse paradigm by initializing 3D reference points and projecting them into all available modalities to sample features. Although methods have recently achieved good performance, there remains notable performance gap compared to dense counterparts.
@@ 3.1 Perspective-Aware Query Generation
Figure 2: Motivations and details of our proposed PAQG module. (a) 3D detectors struggle with low sensitivity when detecting distant and small objects. (b) 2D detectors demonstrate excellent pixel-wise perception capabilities on such objects. (c) the PAQG module adopts the coupled 2D and monocular-3D sub-networks to predict dense boxes under the supervision of a perspective loss. We pick top-ranked boxes to propose high-quality queries, and then interact with camera features via a cross-attention module.
The lightweight perspective detector in the PAQG module consists of the coupled 2D (e.g. FCOS [61]) and monocular-3D (e.g. FCOS3D [66]) sub-networks. Taken the multi-view/scale image features $X_{\text{cam}}$ as input, the monocular-3D sub-network predicts raw 3D attributes, i.e., depths $\mathbf{d}$, rotation angles, sizes, and velocities throughout different views. Simultaneously, the 2D sub-network predicts corresponding 2D attributes, i.e., center coordinates $[\mathbf{c_{x}},\mathbf{c_{y}}]$, confidence scores, and category labels. For each view $v$, we project the box centers into 3D space based on corresponding camera extrinsic ${E_{v}}$ and intrinsic ${I_{v}}$, i.e.,
@@
The 3D center $\mathbf{c}^{3D}$ will combine with the predicted size, rotation angle, and velocity to form 3D boxes. Then, we perform non-maximum suppression in 3D space to filter intersecting boxes and pick the top $N_{k}$ boxes ranked by confidence scores, to initialize queries with image features interacted via an efficient cross-attention module. Formally,
@@
where $\mathcal{P}_{\text{cam}}^{v}(c^{3D}_{i})$ projects the 3D center $c^{3D}_{i}$ to $v$-th image using corresponding camera parameters. Besides, $\mathcal{V}$ denotes the set of hit views. $\mathcal{BS(\cdot)}$ denotes the bilinear sampling function. Since some objects may be overlooked, we preserve $N_{r}$ randomly initialized query boxes. Finally, our PAQG module generates total $N_{q}=N_{k}+N_{r}$ query proposals. This way, our PAQG module provides input-dependent query proposals to elevate the understanding of comprehensive perspective priors (2D and 3D attributes) for 3D detectors, thereby aiding in detecting distant and small objects.
@@ 3.2.3 Channel-Spatial Correlation Aware Mixing
where $W_{s}$ is the spatial correlation shared across different channels. After channel-spatial correlation aware mixing, the features are flattened and aggregated by a linear layer.
@@ 3.3 Uncertainty-Aware Fusion
$\bar{Q}=f_{UA}(F_{\text{cam}},U_{\text{cam}},F_{\text{lid}},U_{\text{lid}}),$ (9)
where $\bar{Q}=\{\bar{q}_{i}\}_{i=1}^{N_{q}}\subset\mathbb{R}^{C}$ and $f_{UA}$ are the refined query feature and uncertainty-aware fusion function, respectively. Besides, $U_{\text{cam}}$ and $U_{\text{lid}}$ are the uncertainty of two modalities.
@@
where $f_{\text{dist}}$ is the distance predictor consisting of MLPs.
@@
Table 1: Quantitative comparisons of SparseLIF with all state-of-the-art 3D detectors on the nuScenes test benchmark. The notion of modality: Camera (C), LiDAR (L), and Temporal (T). $\dagger$: using external training data; $\ddagger$: using TTA and complex model ensemble (e.g. models with different voxel sizes, BEV sizes, backbones/FPNs/heads); §$\S$: we only use very simple self-model ensemble without TTA for SparseLIF-T.
@@
PAI3D [41] LC 24.5 23.3 30.8 23.3 13.1 71.4 74.2
Lift-Attend-Splat [16] LC 24.3 23.8 34.5 32.8 13.3 75.5 74.9
BEVFusion [44] LC 24.2 22.7 32.0 22.2 13.0 75.0 76.1
DeepInteraction [75] LC 23.5 23.3 32.8 22.6 13.0 75.6 76.3
CMT [71] LC 23.3 22.0 27.1 21.2 12.7 75.3 77.0
BEVFusion4D [4] LCT 22.9 22.9 30.2 22.5 13.5 76.8 77.2
EA-LSS [20] LC 23.4 22.8 27.8 20.4 12.4 76.6 77.6
SparseLIF-T§ LCT 24.3 23.1 28.4 15.2 11.7 75.9 77.7
@@
#1 LC V2-99 VoxelNet 66.2 69.0
#2 LC V2-99 VoxelNet 69.8 73.3
#3 LC V2-99 VoxelNet 71.0 74.3
#4 LC V2-99 VoxelNet 70.5 74.1
#5 LC V2-99 VoxelNet 71.2 74.6
#6 LC V2-99 VoxelNet 68.0 70.8
Table 4: Performance analysis of our PAQG module on detection distances and small object classes on the nuScenes validation set, based on SparseLIF-S. The $AP$ scores of traffic cone (T.C.) and barrier at $30m$- are missing since corresponding annotations are unavailable.
PAQG 00-$10m$ $10$-$20m$ $20$-$30m$ $30m$-
mAP 75.1 74.3 65.5 56.9
mAP 76.2 74.8 66.8 58.5
T.C. 82.5 80.6 69.1 -
T.C. 83.8 82.7 70.4 -
Barrier 72.1 77.5 54.5 -
Barrier 76.4 80.0 63.1 -
@@ 4.1.1 Implementation Details
Each model is trained end-to-end using the AdamW optimizer on eight NVIDIA A100 GPUs with a total batch size of $8$. For fair comparisons, we apply the query-denoising strategy [30], commonly used in sparse detection heads, to address the unstable matching problem. Each model is trained for $24$ epochs with a learning rate of $42e-$ 4.
@@ 4.2 Comparisons with State-of-the-Art 3D Object Detectors
As shown in the top part of Tab. 1, without using any test-time augmentation (TTA) or model ensemble, our SparseLIF-T achieves state-of-the-art single-model performance, reaching $77.0\%$ $NDS$ on the nuScenes test benchmark, significantly outperforming all other 3D detectors by a notable margin. In particular, we outperform the most competitive method FusionFormer [19] by $1.9\%$ $NDS$ without using any external training data. Regarding the test benchmark leaderboard in the bottom part of Tab. 1, many competitive methods [41, 16, 44, 75, 71, 4, 20] adopt very complex model ensemble (e.g. assembling models with different voxel sizes, BEV sizes, backbones/FPNs/heads) and TTA, to strive for top ranking on the test leaderboard. Contrarily, we only use very simple self-model ensemble without TTA for SparseLIF-T (i.e., $0.7\%$ $NDS$ improvement), achieving the best performance of $77.7\%$ $NDS$ and ranking 1st on the test leaderboard by the time of paper submission.
@@
We also conduct latency analysis on the nuScenes dataset. We implement SparseLIF using Pytorch without any acceleration operations. The overall latency of SparseLIF-S is $340ms$ on a single NVIDIA A100 GPU. In detail, the detection head (including the PAQG module, the RIAS module, and the UAF module, etc.) only takes about $40ms$, while the camera and LiDAR backbones take the rest of the time, which demonstrates the efficiency of our detector. We can further speed up our detector by configuring the backbones.
Table 5: Robustness studies of SparseLIF on the nuScenes validation set, under challenging scenarios: LiDAR malfunction, camera malfunction and unsynchronization. Abbreviations: construction vehicle (C.V.), pedestrian (Ped.), and traffic cone (T.C.).
@@ 4.3 Ablation Studies
In Tab. 3, we conduct ablation studies on the nuScenes validation set to evaluate the key components in our multi-modality framework, based on the state-of-the-art single-frame detector SparseLIF-S, which yields highly convincing proofs. The RIAS module plays a vital role in our multi-modality detector. Adopting the PAQG module to generate high-quality query proposals, the $mAP$ and $NDS$ are improved by $1.2\%$ and $1.0\%$ respectively. Adopting the UAF module to conduct multi-modality fusion, the $mAP$ and $NDS$ are improved by $0.7\%$ and $0.8\%$ respectively. When SparseLIF-S assembles all modules, the best performance is reached: $71.2\%$ $mAP$ and $74.6\%$ $NDS$.
@@ 4.4 Robustness Studies
We further visualize the predictions of our SparseLIF under the most challenging scenario of limited LiDAR FOV angle of $120^{\circ}$. As presented in Fig. 4, our SparseLIF precisely detects objects in golden circles with LiDAR input malfunctioned, showing the remarkable robustness of our multi-modality detector attributed to the proposed UAF module.
@@ 5 Conclusion
This paper proposes a high-performance fully sparse detector termed SparseLIF for LiDAR-camera-based 3D object detection. Our SparseLIF achieves state-of-the-art performance by enhancing the awareness of rich representations in two modalities. In particular, SparseLIF consists of (1) the PAQG module, which generates high-quality 3D queries with perspective priors to facilitate the perception of small and distant objects; (2) the RIAS module, which further refines prior queries by RoI feature sampling to embrace the fully sparse paradigm with the capability of low latency and integration of more temporal frames; (3) the UAF module, which quantifies the uncertainty of each modality for multi-modality fusion to enhance robustness against sensor noises. The experimental results demonstrate the superiority of our proposed method over all state-of-the-art 3D object detectors on the nuScenes benchmark. In the future, we will explore applications of SparseLIF on other tasks, such as occupancy prediction.